# TA-Cleaner: A Fine-grained Text Alignment Backdoor Defense Strategy for Multimodal Contrastive Learning

### v2

@@ 3.2 Fine-grained text augmentation
When the original text is just a phrase or can’t be further streamlined, such as “a picture of an apple”, then we only keep the core word “apple” as our positive sub-text.
Negative sub-caption generation After extracting the sub-captions, we perform various forms of replacement operations on them. Since we do not know whether the malicious adversary’s attack target is a specific entity, attribute, relationship, or an entire line of text, our replacements need to consider all elements in the text comprehensively. There are three types of sub-text replacement operations, and the replacement method randomly selects one of them: ① Replace the adjectives of the subject and object. If the adjective is missing, skip this step and perform other types of text replacements. ② Replace the relationship verbs. Similarly, if the relationship verb is missing, choose another replacement method. ③ Replace the nouns of the subject and object.
Since the original training data of pre-trained large models are mostly large-scale, noisy, and uncurated image-text pairs captured from the Internet, their data distribution is rich and complex. Therefore, in addition to using standard lexicons like WordNet Fellbaum (2010) as the data foundation for text augmentation, which is similar to CleanCLIP, we also utilize the text-based large model ChatGPT Wu et al. (2023) to generate a repository of alternative content for the three types of replacements. Each type contains 3000 vocabulary words, and it is required that the vocabulary of each alternative repository be as rich as possible, including commonly used medium- to high-frequency words and low-frequency words from the Internet, ensuring that the feature vectors of negative samples are as loosely distributed as possible in the feature space.
3.3 TA-Cleaner: Pos/Neg sub-texts finetuning
Algorithm 1 TA-Cleaner Finetuning Algorithm
@@ 3.3 TA-Cleaner: Pos/Neg sub-texts finetuning
2:     Random select $K$ image-text pairs from $\mathcal{D}_{ft}$, and generate associated positive and negative sub-texts: $ST^{i}_{p}=G_{p}(T_{i}),ST^{i}_{n}=G_{n}(T_{i})$
@@
4:     $Loss=loss_{TA}=\alpha\cdot loss_{CClip}+\beta\cdot loss_{p-n}$
@@
Previous defense efforts have focused on countering backdoor triggers by augmenting image and text self-supervised learning. However, through our prior analysis, we found that the text self-supervision strength of CleanCLIP is insufficient to withstand triggers carefully optimized in the feature space, unless sacrificing the expression capability of clean samples. Therefore, building upon this, we reinforced the text augmentation method by finely optimizing the feature vectors of text through alternating optimization between self-supervised learning and positive-negative sample adversarial learning, enhancing CLIP’s robustness against image backdoor triggers.
@@
In the game-based finetuning between positive and negative samples, we do not perform fine-grained augmentation on all texts, as this would disrupt the alignment of a large number of clean images and texts, thereby reducing the downstream zero-shot accuracy of clean samples. We randomly select $K$ samples from all text data for fine-grained augmentation, obtaining K augmented data, denoted as $\{I_{i},T_{i},ST^{i}_{p},ST^{i}_{n}\}^{K}_{i=1}$. This random selection approach maximally retains the original feature expression capability’s generalization on downstream tasks while achieving our defense objectives. The mapping of these K data points in the feature space is denoted as $\{z^{I}_{i},z^{T}_{i},z^{p}_{i},z^{n}_{i}\}^{K}_{i=1}$.
@@
Our positive-negative sample finetuning loss function consists of two parts: the $loss_{i2t}$ measures the similarity between positive sample images and text and the dissimilarity between negative sample texts and images, thereby minimizing the information difference between positive sample images and text. The $loss_{t2i}$ measures the similarity between positive sample text and images and the dissimilarity between negative sample images and text, thereby minimizing the information difference between positive sample text and images. Both parts jointly optimize the consistency of multi-modal embedding space. The specific mathematical expressions are as follows:
$loss_{i2t}=-\frac{1}{K}\sum_{i=1}^{K}\log\left(\frac{\exp\left(\langle z^{I}_{i},z^{p}_{i}\rangle/{t_{p}}\right)}{\sum_{j=1}^{K}\exp\left(\langle z^{I}_{i},z^{p}_{j}\rangle/{t_{p}}\right)+\sum_{k=1}^{K}\exp\left(\langle z^{I}_{i},z^{n}_{i}\rangle/{t_{n}}\right)}\right),$ (1)
$loss_{t2i}=-\frac{1}{K}\sum_{i=1}^{K}\log\left(\frac{\exp\left(\langle z^{p}_{i},z^{I}_{i}\rangle/{t_{p}}\right)}{\sum_{j=1}^{K}\exp\left(\langle z^{p}_{j},z^{I}_{i}\rangle/{t_{p}}\right)+\sum_{k=1}^{K}\exp\left(\langle z^{n}_{i},z^{I}_{i}\rangle/{t_{n}}\right)}\right),$ (2)
@@
Here, $t_{p}$ and $t_{n}$ are the temperature parameters for positive and negative samples, which control the sensitivity of the loss function to positive and negative samples by adjusting the weight of the similarity score. Specifically, increasing $t_{p}$ enhances the sensitivity of the similarity score of positive samples, leading the loss function to focus more on the differences between positive samples, which may result in less ideal defense effects. Similarly, increasing $t_{n}$ enhances the sensitivity of the similarity score of negative samples, which may lead to excessive learning of negative samples by the model, ignoring the similarity between positive samples and reducing the model’s generalization ability. Therefore, the setting of these two hyper-parameters $t_{p}$ and $t_{n}$ also has a certain degree of influence on the adversarial learning between positive and negative samples. Hence, our total loss function TA-finetuning can be described as follows:
$loss_{TA}=\alpha\cdot loss_{CClip}+\beta\cdot loss_{p-n},$ (4)
where $\alpha$ and $\beta$ are hyper-parameters, representing the weight of $loss_{CClip}$ and $loss_{p-n}$ respectively.
Finally, our complete finetuning steps are given in Algorithm 1.
@@ 4.1 Setup
Dataset and models As a defense technique during the fine-tuning phase, we adopted the fine-tuning setting of Bansal et al. (2023). We utilized the open-source CLIP model from OpenAI Radford et al. (2021a) as the pre-trained clean model, which is trained on a dataset containing 400 million image-text pairs. We selected 500,000 image-text pairs (CC500K) as our fine-tuning dataset from the CC3M dataset Sharma et al. (2018). And following Bansal et al. (2023), we use the ResNet-50 model as the CLIP vision encoder and a transformer as the text encoder during fine-tuning. We conducted our experiments using an A100 GPU.
The victim models generation We also used the CC500K to simulate the adversary’s attack process. Specifically, we randomly selected 1500 samples from CC500K for different kinds of backdoor attacks, embedded the triggers into the images, and modified the corresponding text to attack the target category with a specific template. The other samples are kept unchanged. This dirty dataset is then used to finetune the pre-trained large model. For finetuning, we use a batch size of $128$, an iteration number of $5$, AdamW as our optimizer, a base learning rate of $61e-$ 6, and the number of steps to warm up the learning rate is $10000$, the weight decay of the optimizer is $0.1$. The Adam momentum factor and Adam rmsprop factor are $0.9$ and $0.999$, and the Adam eps is $81e-$ 8.
Defense finetuning We utilize the CC500K to conduct CleanCLIP and our TA-Cleaner finetuning. For both, we use a batch size of $64$, an iteration number of $10$, AdamW as our optimizer. The learning rate warm up steps is $10000$, the weight decay of the optimizer is $0.1$. The Adam momentum factor and Adam rmsprop factor are $0.9$ and $0.999$ respectively, and the Adam eps is $81e-$ 8, the base learning rates are $745e-$ 7 for CleanCLIP and $55e-$ 5 for our TA-Cleaner.
Evaluation metrics Following Yang et al. (2024a); Bansal et al. (2023) and most attacks like Liang et al. (2023), we adopt benign accuracy (BA, $\uparrow$) and attack success rate (ASR, $\downarrow$) as our evaluation metrics. For BA, a higher value indicates superior clean performance, while for ASR, a higher value indicates stronger attack performance. Using these metrics, we assess the defense strategies across two commonly used tasks: zero-shot classification on the ImageNet-1K validation set and linear probe. In the linear probe task, the feature extraction layers remained fixed, and only the linear layer was trained on 50,000 clean images from the ImageNet-1K training set, followed by testing on the ImageNet-1K validation set.
4.2 TA-Cleaner performance
Similar to Bansal et al. (2023) and Yang et al. (2024a), we conduct zero-shot testing on ImageNet1K to evaluate our performance. We utilize six attack methods to generate victim models: BadNet Gu et al. (2017), Blended Chen et al. (2017), SIG Liu et al. (2020), WaNet Nguyen and Tran (2021), SSBA Li et al. (2021b), and BadCLIP Liang et al. (2023). Among them, the first five are classic backdoor attack methods in supervised learning, while BadCLIP is a recently developed attack technique specifically tailored for CLIP. For each attack method, we randomly select 1500 images from CC500K for poisoning and subsequently finetune to generate poisoned models. We apply both CleanCLIP and TA-Cleaner defenses separately to these six poisoned models and obtain the Top-k (k=1,3,5,10) BA (%) and ASR (%) after defense finetuning. Our final results are presented in Table 1. In our implementation of TA-Cleaner, for the first five attack methods, we randomly selected 1,000 images per iteration for positive and negative subtext generation and finetuning. However, for BadCLIP, we randomly sampled 3,000 images for defense, as this is an exceptionally potent attack method where a smaller sample size would be insufficient to generate a defense boundary to resist the proximity of poisoned image features.
@@ 4.2 TA-Cleaner performance
TA-Cleaner 88.91 6.77 88.86 46.72 89.13 15.70 89.12 2.19 90.02 10.27 88.81 28.08
Table 2: The linear-probe classification accuracy (%, $\uparrow$) on a series of datasets.
@@
TA-Cleaner 83.32 60.01 72.37 65.78 95.83 51.06 83.01 77.21 70.35
Regarding Top-1 performance Compared to the victim model (NoDefense), although CleanCLIP already exhibits excellent defense capabilities, our TA-Cleaner can further reduce the ASR, achieving near-zero Top-1 ASR in defending against BadNet and WaNet $(91.28\rightarrow 0.45,91.83\rightarrow 0.15)$. Moreover, for BadCLIP, which CleanCLIP fails to defend against, our method can also significantly weaken its toxicity. This is because BadCLIP’s trigger optimization primarily focuses on adjusting and altering the poisoned image’s feature vector to approach the target text’s features indefinitely, without changing the text’s feature space. In contrast, CleanCLIP’s use of synonym substitution renders the text’s self-supervision insufficient to protect the target text features and fails to sever the backdoor trigger. Our TA-Cleaner method pays more attention to enhancing textual information, achieving fine-grained alignment between random positive and negative subtext features and image features, and actively distancing the target text features to sever the toxic connections, which are attempted by the image triggers.
Regarding Top-k performance In the zero-shot classification test, Top-1 accuracy indicates whether the class with the highest predicted probability by the model matches the true class, focusing on the model’s accuracy in a single prediction. Conversely, Top-k accuracy considers whether the true class is among the top-k predicted classes with the highest probabilities. Even if the class with the highest probability is not the true class, the prediction is deemed correct if the true class is within the model’s top-k predicted classes. Compared to Top-1, Top-k accuracy encompasses a broader prediction scope and thus typically yields higher accuracy. In the Zero-shot classification test, as the model is tasked with handling unseen classes, it may not accurately predict the true class with the highest probability. In such scenarios, Top-k accuracy offers a more comprehensive assessment of the model’s performance, as it takes into account the model’s predictive capability across multiple potentially correct classes. Therefore, we also present the BA and ASR of our defense method on Top-3, Top-5, and Top-10 in Table 1. Compared with NoDefense and CLeanCLIP, it is observed that our TA-Cleaner does not induce overfitting tendencies in the defense processing, as evidenced by the improved defensive performance even within the Top-10 range, while maintaining nearly identical clean sample performance as CleanCLIP. This signifies that our defense strategy does not compromise the usability of the pre-trained model.
@@
About Benign Accuracy Due to our defense is a finetuning-based strategy and the limitation of relevant computational resources, fine-tuning a large pre-trained model with a small dataset (3 million VS. 500K) will inevitably affect the capability of clean feature alignment, i.e., the BA performance. Nevertheless, it is noteworthy that, compared with CleanCLIP, a similarly fine-tuning approach, our proposed TA-Cleaner significantly reduces the ASRs while maintaining almost the same BAs as theirs, which are also shown in Table 1.
The ASR during defense epochs Beyond the ultimate defense consequence, our TA-Cleaner achieves faster and better defense performance compared to CleanCLIP, as illustrated in Figure 3. We present in Figure 3 the ASR trends of CleanCLIP (CClip) and TA-Cleaner with increasing epochs for BadNet, SIG, and BadCLIP attacks. It is observable that TA-Cleaner significantly reduces ASR often within the first epoch and converges relatively steadily thereafter. This implies that merely with the cost of fine-tuning a few thousand additional samples, TA-Cleaner can achieve faster and superior defense performance compared to CleanCLIP.
@@
The availability of TA-Cleaner We evaluated TA-Cleaner using linear-probe methods on a series of datasets introduced by Kornblith et al. (2019) to investigate whether it negatively impacts the model’s usability and transfer performance. For this evaluation, we tested models subjected to BadNet poisoning and defenses, with a learning rate of $31e-$ 3 during linear probe training. The corresponding test results are shown in Table 2. As observed, we achieved test results comparable to CleanCLIP, indicating that we significantly reduced the ASR without compromising the model’s performance and transferability. More details about the dataset information and results are shown in Appendix C.
4.3 Analysis
Ablation As shown in Table 1, Table 2 and Figure 3, when we introduced fine-grained augmentation of positive and negative sub-samples, indicated by the addition of $loss_{p-n}$, our TA-Cleaner significantly improved defense performance compared to the original CleanCLIP, without compromising the model’s performance on clean samples.
@@ 4.3 Analysis
Figure 3: The decline curve of Top-k ASR (%) over epochs of CleanCLIP and our TA-Cleaner, on different backdoor attacks.
@@
(c) TA-texts nums
Figure 4: (a) The Top-1 ASR (%) and BA (%) with different $loss_{p-n}$ weight $\beta/\alpha$. (b) The Top-1 ASR (%) over epochs with different pos/neg temperature targeting BadNet attack. (c) The Top-1 ASR (%) of BadCLIP over epochs with different TA-texts numbers (the num of texts we apply TA-Cleaner in each epoch).

### v3

@@ 3.2 Counterfactual Text Augmentation
When the original text is too simple or contains minimal content (e.g., “a picture of an apple”), we preserve only the most critical word, such as "apple", as the positive sub-caption to retain the core meaning.
Counterfactual: Negative sub-caption generation For each sub-caption, we perform random semantic replacement operations aimed at disrupting the binding between the trigger and the target text. Since we cannot precisely know which entity, attribute, or relationship the adversary might exploit, our replacements are comprehensive and cover all possible elements. We apply three types of replacement operations, and one is selected randomly for each sample: ① Replace the adjectives associated with the subject and object. If no adjectives are present, this step is skipped. ② Replace the relational verbs. If missing, another replacement method is chosen. ③ Replace the subject and object nouns.
Considering the large, noisy, and uncurated nature of pre-trained models’ training data, which captures a rich and diverse data distribution, we augment text using a combination of WordNet [10] and the large language model ChatGPT [33]. The latter helps generate a diverse repository of alternative words, ensuring that negative subtexts are loosely distributed in feature space. Each replacement repository contains 3000 terms, ranging from common to rare words, ensuring maximum disruption of the backdoor trigger’s semantic binding.
3.3 CleanerCLIP: Fine-grained Counterfactual Semantic Finetuning
Algorithm 1 CleanerCLIP Finetuning Algorithm
@@ 3.3 CleanerCLIP: Fine-grained Counterfactual Semantic Finetuning
2:     Random select $K$ image-text pairs from $\mathcal{D}_{ft}$, and generate associated positive and negative sub-captions: $ST^{i}_{p}=G_{p}(T_{i}),ST^{i}_{n}=G_{n}(T_{i})$
@@
4:     $Loss=loss_{Cleaner}=\alpha\cdot loss_{CClip}+\beta\cdot loss_{p-n}$
@@
Previous defense strategies have largely concentrated on leveraging image and text self-supervised learning to counter backdoor triggers. However, our analysis reveals a critical limitation in CleanCLIP: its self-supervision strength is insufficient to defend against meticulously crafted triggers that exploit existing clean features without introducing additional toxic embedding clusters. Specifically, while CleanCLIP reinforces clean feature representations, it fails to address subtle manipulations of these features by attackers who strategically leverage the clean features themselves. This gap highlights the need for a more robust defense against advanced, stealthy attack techniques during finetuning. To overcome this limitation, we propose a fine-grained text semantic augmentation approach that incorporates both positive and negative sub-captions during the fine-tuning process. This method optimizes text features by employing alternating optimization between self-supervised learning and adversarial sample generation. By reinforcing the robustness of text representations without compromising the expressiveness of clean samples, this approach provides enhanced protection against image-based backdoor triggers, aligning with the lightweight defense mechanism we advocate for.
@@
We avoid fine-grained semantic augmentation on the whole dataset during fine-tuning to prevent disrupting the alignment of clean images and texts, which may lower the downstream zero-shot accuracy. We randomly select $K$ samples from all text data for fine-grained augmentation, obtaining $K$ augmented data, denoted as $\{I_{i},T_{i},ST^{i}_{p},ST^{i}_{n}\}^{K}_{i=1}$. This random selection approach maximally retains the original feature expression capability’s generalization on downstream tasks while achieving our defense objectives, and effectively minimizing the additional computational cost of the fine-tuning defense. The mapping of these $K$ data points in the feature space is denoted as $\{z^{I}_{i},z^{T}_{i},z^{p}_{i},z^{n}_{i}\}^{K}_{i=1}$. Note that here, $z^{p}_{i}$ and $z^{n}_{i}$ represent the feature embeddings of positive and negative subtexts, respectively.
@@
Our factual-counterfactual finetuning loss function consists of two parts: the $loss_{i2t}$ measures the similarity between positive sample images and text and the dissimilarity between negative sample texts and images, thereby minimizing the information difference between positive sample images and text. The $loss_{t2i}$ measures the similarity between positive sample text and images and the dissimilarity between negative sample images and text, thereby minimizing the information difference between positive sample text and images. Both parts jointly optimize the consistency of multi-modal embedding space. The specific mathematical expressions are as follows:
$loss_{i2t}=-\frac{1}{K}\sum_{i=1}^{K}\log\left(\frac{\exp\left(\langle z^{I}_{i},z^{p}_{i}\rangle/{t_{p}}\right)}{\sum_{j=1}^{K}\exp\left(\langle z^{I}_{i},z^{p}_{j}\rangle/{t_{p}}\right)+\sum_{k=1}^{K}\exp\left(\langle z^{I}_{i},z^{n}_{i}\rangle/{t_{n}}\right)}\right)$
$loss_{t2i}=-\frac{1}{K}\sum_{i=1}^{K}\log\left(\frac{\exp\left(\langle z^{p}_{i},z^{I}_{i}\rangle/{t_{p}}\right)}{\sum_{j=1}^{K}\exp\left(\langle z^{p}_{j},z^{I}_{i}\rangle/{t_{p}}\right)+\sum_{k=1}^{K}\exp\left(\langle z^{n}_{i},z^{I}_{i}\rangle/{t_{n}}\right)}\right)$
@@
Here, $t_{p}$ and $t_{n}$ are the temperature parameters for positive and negative samples, which control the sensitivity of the loss function to positive and negative samples by adjusting the weight of the similarity score. Specifically, increasing $t_{p}$ enhances the sensitivity of the similarity score of positive samples, leading the loss function to focus more on the differences between positive samples, which may result in less ideal defense effects. Similarly, increasing $t_{n}$ enhances the sensitivity of the similarity score of negative samples, which may lead to excessive learning of negative samples by the model, ignoring the similarity between positive samples and reducing the model’s generalization ability. Therefore, the setting of these two hyper-parameters $t_{p}$ and $t_{n}$ also has a certain degree of influence on the adversarial learning between positive and negative samples. Hence, our total loss function $loss_{Cleaner}$ can be described as follows:
$loss_{Cleaner}=\alpha\cdot loss_{CClip}+\beta\cdot loss_{p-n},$ (4)
where $\alpha$ and $\beta$ are hyper-parameters, representing the weight of $loss_{CClip}$ and $loss_{p-n}$ respectively. Finally, our complete finetuning steps are given in Algorithm 1.
@@ 4.1 Setup
Dataset and models As a defense technique during the fine-tuning phase, we adopted the fine-tuning setting of [1]. We utilized the open-source CLIP model from OpenAI [25] as the pre-trained clean model, which is trained on a dataset containing 400 million image-text pairs. We selected 500,000 image-text pairs (CC500K) as our fine-tuning dataset from the CC3M dataset [28]. Following [1], we use the ResNet-50 model as the CLIP vision encoder and a transformer as the text encoder during fine-tuning. We conducted our experiments using an A100 GPU.
@@
CleanerCLIP 88.91 6.77 88.86 46.72 89.13 15.70 89.12 2.19 90.02 10.27 88.81 28.08
The victim models generation We utilized the CC500K dataset to simulate the adversary’s attack process. Specifically, we randomly selected 1,500 samples from CC500K for various types of backdoor attacks, embedding triggers into the images. The corresponding text was modified to target specific categories using a predefined template, while the remaining samples were kept unchanged. This contaminated dataset was then used to fine-tune the pre-trained CLIP model. For fine-tuning, we employed a batch size of 128, an iteration count of 5, and a base learning rate of $1\times 10^{-6}$. The learning rate was warmed up over 10,000 steps, using AdamW as our optimizer with a weight decay of 0.1. The Adam momentum factor and RMSProp factor were set to 0.9 and 0.999, respectively, with an epsilon value of $1\times 10^{-8}$. And the attack target label is “banana".
Defense finetuning We employed the CC500K dataset to conduct clean fine-tuning (FT), CleanCLIP, and our proposed CleanerCLIP. For both methods, we utilized a batch size of 64, an iteration count of 10, and AdamW as the optimizer. The learning rate was warmed up over 10,000 steps, with a weight decay of 0.1 for the optimizer. The Adam momentum factor and RMSProp factor were set to 0.9 and 0.999, respectively, with an epsilon value of $1\times 10^{-8}$. The base learning rate for both methods was set to $4.5\times 10^{-6}$.
Evaluation metrics Following [37, 1] and most attacks like [21], we adopt benign accuracy (BA, $\uparrow$) and attack success rate (ASR, $\downarrow$) as our evaluation metrics. For BA, a higher value indicates superior clean performance, while for ASR, a lower value reflects better defense performance. These metrics are used to assess defense strategies across two common tasks: zero-shot classification on the ImageNet-1K validation set and linear probing. In the linear probing test, the feature extraction layers remain fixed, and only the linear layer is trained on clean images from the training set, followed by testing on the validation set.
To comprehensively assess the impact of our defense on CLIP performance, we use Top-k ($k=1,3,5,10$) evaluations for both BA and ASR. Top-k accuracy considers not only the highest probability class predicted by the model but also other high-probability classes, thus offering a more accurate reflection of the model’s generalization capacity in multi-class tasks. This consideration has been widely used in prior works on the CLIP community, such as EVA-CLIP [30], CosmoCLIP [15], and CEIA [34]. However, to our knowledge, previous backdoor defense works targeting CLIP have not reported Top-k metrics, leading to incomplete performance evaluation. Therefore, by incorporating Top-k BA/ASR evaluations, our work not only provides a more thorough investigation into CLIP’s backdoor vulnerabilities but also demonstrates state-of-the-art performance by significantly reducing Top-k ASR.
@@
CleanerCLIP 83.32 60.01 72.37 65.78 95.83 51.06 83.01 77.21 70.35
4.2 CleanerCLIP performance
Similar to [1] and [37] of pretraining defense methods, we conduct zero-shot testing on ImageNet1K to evaluate our performance. We utilize six attack methods to generate victim models: BadNet [13], Blended [8], SIG [22], WaNet [23], SSBA [20], and BadCLIP [21]. Among them, the first five are classic backdoor attack methods in supervised learning, while BadCLIP is a recently developed attack technique specifically tailored for CLIP. For each attack method, we randomly select 1500 images from CC500K for poisoning and subsequently finetune to generate poisoned models. We apply FT, CleanCLIP, and CleanerCLIP defenses separately to these six poisoned models and obtain the Top-k (k=1,3,5,10) BA (%) and ASR (%) after defense finetuning. Our final results are presented in Table 1. In our implementation of CleanerCLIP, for the first five attack methods, we randomly selected 1,000 images per iteration for positive and negative subtext generation and finetuning. However, for BadCLIP, we randomly sampled 3,000 images for defense, as this is an exceptionally potent attack method where a smaller sample size would be insufficient to generate a defense boundary to resist the proximity of poisoned image features.
Regarding Top-1 performance Compared to the victim model (NoDefense), fine-tuning (FT) improves defense, but certain vulnerabilities remain. For instance, the Top-1 ASR for the BadNet attack is still high at 62.14%, reflecting limitations in effectively countering attacks. CleanCLIP shows stronger defensive capabilities, but our CleanerCLIP reduces ASR even further, achieving near-zero Top-1 ASR against both BadNet and WaNet (from 91.28% to 0.45% and from 91.83% to 0.15%, respectively). Against BadCLIP, which poses significant challenges for CleanCLIP, our method also markedly weakens its effectiveness. BadCLIP’s optimization aligns poisoned images closely with target text features, creating a strong trigger association. CleanCLIP’s EDA-based augmentation does not alter semantic content sufficiently to disrupt this binding, allowing triggers to persist within clean features. Our CleanerCLIP, however, uses random counterfactual semantic enhancement to disrupt trigger connections, while factual positive subtexts strengthen defense by preserving clean sample features during fine-tuning.
Regarding Top-k performance In zero-shot classification, Top-1 accuracy reflects the model’s success in identifying the correct class as the top prediction, while Top-k accuracy assesses whether the true class is among the top-k predictions, thus offering a broader evaluation. Given the inherent challenges in zero-shot tasks, Top-k accuracy often provides a more comprehensive measure of performance, as it takes into account the model’s predictive capability across multiple potentially correct classes. We report BA and ASR for Top-3, Top-5, and Top-10 in Table 1. Unlike NoDefense, FT, and CleanCLIP, our CleanerCLIP avoids overfitting and shows improved defense even across broader prediction scopes. Notably, when FT and CleanCLIP almost fail against BadCLIP (Top-10 ASR at 99.37% and 91.96%, respectively), our method significantly lowers ASR by 71.75% relative to the original victim model, underscoring the robustness of our approach.
@@ 4.2 CleanerCLIP performance
About BA Due to our defense being a finetuning-based strategy and the limitation of relevant computational resources, fine-tuning a large pre-trained model with a small dataset (3 million VS. 500K) will inevitably affect the capability of clean feature alignment, i.e., the BA performance. Nevertheless, it is noteworthy that, compared with CleanCLIP, a similarly fine-tuning approach, our proposed CleanerCLIP significantly reduces the ASRs while maintaining almost the same BAs as theirs, as shown in Table 1.
@@
The availability of CleanerCLIP We evaluated CleanerCLIP using linear-probe methods on a series of datasets introduced by [18] to investigate whether it negatively impacts the model’s usability and transfer performance. For this evaluation, we tested models subjected to BadNet poisoning and defenses, with a learning rate of $31e-$ 3 during linear probe training. The corresponding test results are shown in Table 2. As observed, we achieved test results comparable to CleanCLIP, indicating that we significantly reduced the ASR without compromising the model’s performance and transferability. More details about the dataset information are shown in Sec 8 in the supplementary material.
@@
Figure 3: The decline curve of Top-k ASR (%) over epochs of CleanCLIP and our CleanerCLIP, on different backdoor attacks.
@@
(c) Neg-sample nums
Figure 4: (a) The Top-1 ASR (%) and BA (%) with different $loss_{p-n}$ weight $\beta/\alpha$. (b) The Top-1 ASR (%) over epochs with different pos/neg temperature targeting BadNet attack. (c) The Top-1 ASR (%) of BadCLIP over epochs with different Neg-sample numbers (the number of texts we apply CleanerCLIP in each epoch).
The ASR during defense epochs Beyond the ultimate defense consequence, our CleanerCLIP achieves faster and better defense performance compared to CleanCLIP, as illustrated in Figure 3. We present in Figure 3 the ASR trends of CleanCLIP (CClip) and CleanerCLIP with increasing epochs for BadNet, SIG, and BadCLIP attacks. It is observable that CleanerCLIP significantly reduces ASR often within the first epoch and converges relatively steadily thereafter. This implies that merely with the cost of fine-tuning a few thousand additional samples, CleanerCLIP can achieve faster and superior defense performance compared to CleanCLIP.
4.3 Analysis
Ablation As shown in Table 1, Table 2 and Figure 3, when we introduced fine-grained augmentation of positive and negative sub-samples, indicated by the addition of $loss_{p-n}$, our CleanerCLIP significantly improved defense performance compared to the original CleanCLIP, without compromising the model’s performance on clean samples.